OpenStudy (anonymous):

Given that 2sintheta=3costheta, find tan theta

OpenStudy (shubhamsrg):

tan theta = sintheta / cos theta

OpenStudy (anonymous):

We know that \[\tan \theta = \frac{ \sin \theta }{ \cos \theta }\] Try rearranging your equation to get sin on top of cos

OpenStudy (anonymous):

would one of them become negative then?

OpenStudy (anonymous):

\[2 \sin \theta = 3 \cos \theta\]Divide by cos\[\frac{ 2 \sin \theta }{ \cos \theta } = 3\]Divide by 2\[\frac{ \sin \theta }{ \cos \theta } = 3/2\] With me so far?

OpenStudy (anonymous):